# PainFormer: a Vision Foundation Model for Automatic Pain Assessment

### v2

@@ Abstract
Pain is a manifold condition that impacts a significant percentage of the population. Accurate and reliable pain evaluation for the people suffering is crucial to developing effective and advanced pain management protocols. Automatic pain assessment systems provide continuous monitoring and support decision-making processes, ultimately aiming to alleviate distress and prevent functionality decline. This study introduces PainFormer, a vision foundation model based on multi-task learning principles trained simultaneously on 14 tasks/datasets with a total of 10.9 million samples. Functioning as an embedding extractor for various input modalities, the foundation model provides feature representations to the Embedding-Mixer, a transformer-based module that performs the final pain assessment. Extensive experiments employing behavioral modalities–including RGB, synthetic thermal, and estimated depth videos–and physiological modalities such as ECG, EMG, GSR, and fNIRS revealed that PainFormer effectively extracts high-quality embeddings from diverse input modalities. The proposed framework is evaluated on two pain datasets, BioVid and AI4Pain, and directly compared to 75 different methodologies documented in the literature. Experiments conducted in unimodal and multimodal settings demonstrate state-of-the-art performances across modalities and pave the way toward general-purpose models for automatic pain assessment. The foundation model’s architecture (code) and weights will be made available after the paper’s final publication.
@@ I Introduction
Pain is a pervasive medical issue worldwide, impacting up to $30\%$ of the general adult population [2] and between $83\%$ and $93\%$ of senior adults living in residential care facilities [7]. According to the Global Burden of Disease (GBD) study, pain is the leading cause of years lived with disability (YLD) [8], with three of the top contributors being chronic pain conditions: back pain, musculoskeletal disorders, and neck pain [9]. The influence of pain expands beyond individuals to affect society, presenting clinical, economic, and social challenges. The estimated economic and healthcare costs associated with pain due to reduced work productivity in the United States range from \$$\$560$ to \$$\$635$ billion, exceeding the costs of heart disease, cancer, and diabetes combined [10]. Europe shows similar patterns, with direct healthcare expenses and indirect socioeconomic costs of chronic pain accounting for $3\%$ to $10\%$ of the gross domestic product [11]. Additionally, in Australia, the average annual cost for one of the $15.4\%$ of residents living with chronic pain is between AU\$$\$22,588$ and AU\$$\$42,979$ when including non-financial costs [12]. Pain extends beyond direct impacts on a patient’s life, leading to a range of adverse effects, including opioid use, drug overuse, addiction, deteriorating social relationships, and psychological disorders [13]. Over the past two decades, the use of prescription opioids has significantly increased in the United States, where the rate of overdose deaths has more than quadrupled from $1999$ to $2016$ [14]. Additionally, the side effects of these opioids, including lethargy, depression, anxiety, and nausea, significantly affect workforce productivity and overall quality of life [15].
@@
Accurate pain assessment is vital for early diagnosis, disease progression monitoring, and treatment effectiveness evaluation, especially in managing chronic pain [16]. This importance has led to pain being designated as “the fifth vital sign” in nursing literature [17]. Furthermore, pain evaluation is essential in physiotherapy, where the therapist externally induces stimuli, and understanding the patient’s pain level is needed [18]. Objectively evaluating pain is crucial for delivering appropriate care, particularly for vulnerable groups who cannot directly or reliably communicate their pain condition, such as infants, young children, those with mental health conditions, and elderly individuals. Various methodologies are employed to assess pain, such as self-reporting, where individuals describe their pain experiences, which is currently regarded as the gold standard [19]. Pain evaluation methods in clinical settings offer quantifiable indicators of pain, ranging from the Numeric Pain Rating Scale (NPRS) and Visual Analogue Scale (VAS) to quantitative sensory testing methods such as the pressure pain detection threshold (PPDT) [20]. Additionally, behavioral indicators such as facial expressions (like grimacing, open mouth, or lifted eyebrows), vocalizations (such as crying, moaning, or screaming), and body and head movements are important markers [21]. Physiological measurements, including electrocardiography (ECG), electromyography (EMG), galvanic skin responses (GSR), and respiration rate, also provide essential understandings of the physiological manifestations of pain [16]. Furthermore, brain monitoring techniques such as near-infrared spectroscopy (fNIRS) have shown the capability to detect hemodynamic activity changes related to pain stimuli [22].
@@
Over the past two decades, computational science researchers have dedicated their efforts to developing models and algorithms for advancing automatic pain recognition systems [23]. The aim is to accurately identify the presence and intensity of pain by analyzing physiological and behavioral responses. The emergence of deep learning and artificial intelligence (AI) methods has further explored these automatic approaches, aiming at interpreting the complex, multifaceted nature of pain [16]. Multiple studies highlight the potential of automated systems utilizing behavioral or physiological pain assessment modalities [24]. Sario et al. [25] demonstrate the viability of accurately recognizing pain via facial expressions, illustrating their value in clinical settings. Multimodal sensing appears especially promising, exhibiting increased accuracy in pain monitoring systems [21]. Additionally, incorporating the temporal dimension of the modalities has been demonstrated to enhance the effectiveness of pain assessment [16]. In recent years, affective computing research has increasingly adopted thermal imaging techniques [26], motivated by literature indicating that stress and cognitive load have notable effects on skin temperature [27]. These effects arise from the autonomic nervous system’s (ANS) control over physiological signals such as heart rate, respiration rate, blood perfusion, and body temperature, all of which reflect human emotions and affects [26]. Moreover, muscle contractions can affect facial temperature by transferring heat to the facial skin [28]. Thus, thermal imaging has emerged as a promising technique for capturing transient facial temperature [29]. However, only a few studies have explored thermal imaging in pain research. In [30], researchers reported that facial temperature increases after exposure to a painful stimulus, suggesting that thermal cameras might be effective tools for monitoring pain. In [31], the authors demonstrated that thermal videos achieved similar accuracy to RGB videos in recognizing pain in an automatic pain assessment environment. Moreover, our previous study [32] introduced synthetic thermal videos through a deep-learning generative process and evaluated their effectiveness in recognizing pain. The findings showcased that the performances of the synthetic modality are equivalent to the original RGB videos, while the combination of them holds significant potential. Furthermore, depth cameras are employed in emotion recognition research because they enable the extraction and use of features related to head pose and 3D facial landmark analysis [33]. These camera sensors capture body and head dynamics, facilitating the analysis of pixel and depth intensity changes associated with movements and affects [34]. In addition, depth sensors are more invariant and maintain their effectiveness under poor or uneven lighting conditions [35]. The research in [36] demonstrated that depth information combined with RGB enhanced the accuracy of micro-expression analysis and, by extension, emotion recognition tasks, while in [37], the authors utilized the 3D positions of facial feature points to identify basic emotions. In pain research, the adoption of depth modality is limited. The authors in [38] utilized depth map videos to estimate head movements, which enhanced pain recognition performance when combined with facial expressions.
@@ II Related work
Various compelling methods have been introduced, utilizing video modalities, ranging from hand-crafted feature engineering for facial expressions analysis to employing raw videos with sophisticated deep learning architectures. Werner et al. [47] analyzed facial expressions by extracting features based on point distances and head poses, which were estimated using depth information. By integrating these into a unified feature vector and applying a random forest classifier, they achieved a $76.60\%$ accuracy rate in the binary pain classification task. Patania et al. [48] calculated fiducial points and applied graph neural networks (GNN) to achieve $73.20\%$ accuracy, while the authors in [49] employed vision transformers on raw videos to reach an accuracy of $77.10\%$. Other researchers have made efforts to compute biosignal-related information directly from videos, aiming to enhance assessment accuracy while minimizing the reliance on contact sensors. Yang et al. [50] leveraged raw videos and 3D CNNs to extract facial features and estimate remote photoplethysmography, effectively combining behavioral and physiological cues for enhanced performance. Similarly, the authors in [51] used 3D CNNs to extract pseudo-heart rate from videos, achieving over a $10\%$ increase in accuracy by integrating facial features with heart rate data compared to using videos in isolation.
@@ III Methodology
This section describes the architecture of the proposed framework’s components. It details the foundation model’s multitask learning-based pretraining, augmentation techniques, and training settings for pretraining and pain assessment tasks. Additionally, it explains the generation of synthetic thermal videos, depth estimation for creating depth videos, and visualization of biosignal modalities used in this study.
@@ III-A1 PainFormer
Before discussing the specific components, we provide preliminary information to outline the general concept of the PainFormer architecture. Every 2D input image $I$ is divided into $n$ non-overlapping patches, with each patch $\in\mathbb{R}^{n\times h\times w\times 3}$, where $h$ and $w$ define the resolution of each patch and are equal to $16\times 16$, and the $3$ indicates the number of color channels. Each patch is linearly projected into a token with dimension $d=768$, followed by a positional encoding layer. Applying Discrete Fourier Transform (DFT), to a 1D input sequence of $N$ elements, $x[n]$, where $n$ ranges from 00 to $N-1$, results in the following expression:
$X[k]=\sum_{n=0}^{N-1}x[n]\cdot e^{-i2\pi\tfrac{k}{N}n}:=\sum_{n=0}^{N-1}x[n]\cdot W_{N}^{kn},$ (1)
where $i$ denotes the imaginary unit, and $W$ represents the twiddle factor, defined as $W_{N}=e^{-i\tfrac{2\pi}{N}}$. In this manner, the sequence is transformed into the frequency domain. Conversely, the original input sequence can be reconstructed by applying the inverse Discrete Fourier Transform (IDFT):
@@
where $x[n]$ is the initial time-domain sequence. Furthermore, Eq. (1) can be adapted for two-dimensional inputs, $x[m,n]$, with $10\leq m\leq M-$ 1 and $10\leq n\leq N-$ 1, characterized by:
@@
where DWConv denotes a depthwise convolution layer. In addition, layer normalization is employed before and after the FFT and IFFT processes, refer to Fig. 1(b).
@@ III-A3 Video-Encoder
The architecture of the specific module resembles the Embedding-Mixer; however, for efficiency, it comprises just $1$ layer featuring a $1$ cross-attention module with $1$ head. The number of latent variables, $n$, is set to $256$, while the output embedding length is $40$. This module is employed only within the particular framework for one of the multimodal approaches presented, where video embeddings are integrated with GSR embeddings. The module is illustrated in Fig. 1(e).
@@ III-C Biosignal Visualization
Since the foundation model presented in this study is vision-based, a 2D representation of physiological modalities needs to be employed. Four different visualizations are explored: (1) waveform diagrams, which depict the shape and form of a signal as it progresses over time, representing amplitude, frequency, and phase; (2) spectrogram-angle, which illustrates the phase angles of the frequencies; (3) spectrogram-phase, which shows phase information with unwrapping to address discontinuities; and (4) spectrogram-PSD, which represents the power spectral density, exhibiting how power is distributed across frequencies over time. Fig. 3 presents an example of each of these four visualizations.
@@ III-D Foundation Training
TABLE III: Datasets utilized for the multitask learning-based pretraining process of the PainFormer.
@@ III-F1 BioVid Heat Pain Database
The particular dataset is widely known and established in the field of pain research. It includes facial videos, electrocardiograms, electromyograms, and galvanic skin response measurements from eighty-seven $(n\text{=}87)$ healthy individuals ($44$ males and $43$ females, aged between $20$ and $65$). The experimental setup involved using a thermode to induce pain in the participants’ right arm. Before data collection, each participant’s pain and tolerance thresholds were determined. These thresholds marked the minimum and maximum levels of pain, with two additional intermediate levels, leading to a total of five distinct pain intensities: No Pain (NP), Mild Pain (P1), Moderate Pain (P2), Severe Pain (P3), and Very Severe Pain (P4). The temperature settings used for inducing pain varied within the range from P1 to P4 but never exceeded $50.5^{\circ}\text{C}$. Each participant experienced $20$ pain inductions at each of the four predetermined levels of intensity (P1 to P4). Each stimulus lasted $4s$, followed by a recovery period of $8$ to $12s$. Additionally, $20$ baseline measurements were taken at $32^{\circ}\text{C}$ (NP), resulting in $100$ stimulations per participant, delivered randomly. The data was then processed to create segments of $5.5s$ starting $1s$ after reaching the target temperature for each stimulation. This processing yielded $8,700$ samples, each lasting $5.5s$, distributed evenly across the five pain intensity classes for each modality, covering all $87$ subjects. The videos are recorded at a frame rate of $25$ frames per second (FPS), while the biosignal (ECG, EMG, GSR) recordings are sampled at a rate of 515151512 Hz.
@@ III-F2 AI4Pain Dataset
The AI4Pain Grand Challenge 2024 dataset is a more recent addition to the field and is specifically designed for advanced pain recognition tasks using fNIRS and facial video data. Sixty-five $(n\text{=}65)$ volunteers, including $23$ females, took part in the experiment, with their ages ranging from $17$ to $52$ years. While the dataset also includes physiological signals such as photoplethysmography (PPG), electrodermal activity (EDA), and respiration (RESP), these additional data are not yet publicly available. The current version of the AI4Pain Challenge dataset is divided into three parts: training ($41$ Volunteers), validation ($12$ Volunteers), and testing ($12$ Volunteers). The fNIRS and video recording setup used in this dataset provides comprehensive data on both brain activity and facial movements. The fNIRS data was recorded using an Artinis device (Artinis Medical Systems, Gelderland, the Netherlands). This device measures changes in the concentrations of oxygenated haemoglobin (HBO2) and deoxygenated haemoglobin (HHB) (in $\mu$mol/L). The fNIRS system includes $24$ channels covering the prefrontal cortex, with optodes ($10$ sources and $8$ detectors) separated by $30$ mm. Near-infrared light is emitted by sources with wavelengths of $760$ nm and $840$ nm, with a sampling rate of $50$ Hz. The second sensing technology is a video camera (Logitech StreamCam) that captures facial video data and movements at a sampling rate of $30$ FPS. The AI4Pain dataset is stratified into three distinct categories of varying levels of pain intensity: No Pain, Low Pain, and High Pain. Specifically, the dataset consists of $65$ instances (each lasting $60s$) of No Pain, $780$ instances of Low Pain (each lasting $10$s), and $780$ instances of High Pain (each lasting $10$s). The No Pain category includes instances from the baseline period at the start of each experiment, providing fNIRS and facial video data for comparison with pain-induced responses. The Low Pain category comprises instances of mild pain based on the pain tolerance test, capturing subtle neurological and behavioural changes in the corresponding fNIRS and video data. Finally, the High Pain category consists of instances where subjects experienced significant pain, also based on the pain tolerance test, leading to notable physiological and behavioural responses in the fNIRS and video data.

### v3

@@ Abstract
Pain is a manifold condition that impacts a significant percentage of the population. Accurate and reliable pain evaluation for the people suffering is crucial to developing effective and advanced pain management protocols. Automatic pain assessment systems provide continuous monitoring and support decision-making processes, ultimately aiming to alleviate distress and prevent functionality decline. This study introduces PainFormer, a vision foundation model based on multi-task learning principles trained simultaneously on 14 tasks/datasets with a total of 10.9 million samples. Functioning as an embedding extractor for various input modalities, the foundation model provides feature representations to the Embedding-Mixer, a transformer-based module that performs the final pain assessment. Extensive experiments employing behavioral modalities–including RGB, synthetic thermal, and estimated depth videos–and physiological modalities such as ECG, EMG, GSR, and fNIRS revealed that PainFormer effectively extracts high-quality embeddings from diverse input modalities. The proposed framework is evaluated on two pain datasets, BioVid and AI4Pain, and directly compared to 75 different methodologies documented in the literature. Experiments conducted in unimodal and multimodal settings demonstrate state-of-the-art performances across modalities and pave the way toward general-purpose models for automatic pain assessment. The foundation model’s architecture (code) and weights are available at https://github.com/GkikasStefanos/PainFormer.
@@ I Introduction
Pain is a pervasive medical issue worldwide, impacting up to $30\%$ of the general adult population [2] and between $83\%$ and $93\%$ of senior adults living in residential care facilities [7]. According to the Global Burden of Disease (GBD) study, pain is the leading cause of years lived with disability (YLD) [8], with three of the top contributors being chronic pain conditions: back pain, musculoskeletal disorders, and neck pain [9]. The influence of pain expands beyond individuals to affect society, presenting clinical, economic, and social challenges. The estimated economic and healthcare costs associated with pain due to reduced work productivity in the United States range from \$560\\$560 to \$635\\$635 billion, exceeding the costs of heart disease, cancer, and diabetes combined [10]. Europe shows similar patterns, with direct healthcare expenses and indirect socioeconomic costs of chronic pain accounting for $3\%$ to $10\%$ of the gross domestic product [11]. Additionally, in Australia, the average annual cost for one of the $15.4\%$ of residents living with chronic pain is between AU\$22,588\\$22,588 and AU\$42,979\\$42,979 when including non-financial costs [12]. Pain extends beyond direct impacts on a patient’s life, leading to a range of adverse effects, including opioid use, drug overuse, addiction, deteriorating social relationships, and psychological disorders [13]. Over the past two decades, the use of prescription opioids has significantly increased in the United States, where the rate of overdose deaths has more than quadrupled from $1999$ to $2016$ [14]. Additionally, the side effects of these opioids, including lethargy, depression, anxiety, and nausea, significantly affect workforce productivity and overall quality of life [15].
@@
Accurate pain assessment is vital for early diagnosis, disease progression monitoring, and treatment effectiveness evaluation, especially in managing chronic pain [16]. This importance has led to pain being designated as “the fifth vital sign” in nursing literature [17]. Furthermore, pain evaluation is essential in physiotherapy, where the therapist externally induces stimuli, and understanding the patient’s pain level is needed [18]. Objectively evaluating pain is crucial for delivering appropriate care, particularly for vulnerable groups who cannot directly or reliably communicate their pain condition, such as infants, young children, those with mental health conditions, and elderly individuals. Various methodologies are employed to assess pain, such as self-reporting, where individuals describe their pain experiences, which is currently regarded as the gold standard [19]. Pain evaluation methods in clinical settings offer quantifiable indicators of pain, ranging from the Numeric Pain Rating Scale (NPRS) and Visual Analogue Scale (VAS) to quantitative sensory testing methods such as the pressure pain detection threshold (PPDT) [20]. Additionally, behavioral indicators such as facial expressions (like grimacing, open mouth, or lifted eyebrows), vocalizations (such as crying, moaning, or screaming), and body and head movements are important markers [21]. Physiological measurements, including electrocardiography (ECG), electromyography (EMG), galvanic skin responses (GSR), and respiration rate, also provide essential understandings of the physiological manifestations of pain [16, 22]. Furthermore, brain monitoring techniques such as near-infrared spectroscopy (fNIRS) have shown the capability to detect hemodynamic activity changes related to pain stimuli [23, 24].
@@
Over the past two decades, computational science researchers have dedicated their efforts to developing models and algorithms for advancing automatic pain recognition systems [25]. The aim is to accurately identify the presence and intensity of pain by analyzing physiological and behavioral responses. The emergence of deep learning and artificial intelligence (AI) methods has further explored these automatic approaches, aiming at interpreting the complex, multifaceted nature of pain [16]. Multiple studies highlight the potential of automated systems utilizing behavioral or physiological pain assessment modalities [26]. Sario et al. [27] demonstrate the viability of accurately recognizing pain via facial expressions, illustrating their value in clinical settings. Multimodal sensing appears especially promising, exhibiting increased accuracy in pain monitoring systems [21]. Additionally, incorporating the temporal dimension of the modalities has been demonstrated to enhance the effectiveness of pain assessment [16]. In recent years, affective computing research has increasingly adopted thermal imaging techniques [28], motivated by literature indicating that stress and cognitive load have notable effects on skin temperature [29]. These effects arise from the autonomic nervous system’s (ANS) control over physiological signals such as heart rate, respiration rate, blood perfusion, and body temperature, all of which reflect human emotions and affects [28]. Moreover, muscle contractions can affect facial temperature by transferring heat to the facial skin [30]. Thus, thermal imaging has emerged as a promising technique for capturing transient facial temperature [31]. However, only a few studies have explored thermal imaging in pain research. In [32], researchers reported that facial temperature increases after exposure to a painful stimulus, suggesting that thermal cameras might be effective tools for monitoring pain. In [33], the authors demonstrated that thermal videos achieved similar accuracy to RGB videos in recognizing pain in an automatic pain assessment environment. Moreover, our previous study [34] introduced synthetic thermal videos through a deep-learning generative process and evaluated their effectiveness in recognizing pain. The findings showcased that the performances of the synthetic modality are equivalent to the original RGB videos, while the combination of them holds significant potential. Furthermore, depth cameras are employed in emotion recognition research because they enable the extraction and use of features related to head pose and 3D facial landmark analysis [35]. These camera sensors capture body and head dynamics, facilitating the analysis of pixel and depth intensity changes associated with movements and affects [36, 37]. In addition, depth sensors are more invariant and maintain their effectiveness under poor or uneven lighting conditions [38]. The research in [39] demonstrated that depth information combined with RGB enhanced the accuracy of micro-expression analysis and, by extension, emotion recognition tasks, while in [40], the authors utilized the 3D positions of facial feature points to identify basic emotions. In pain research, the adoption of depth modality is limited. The authors in [41] utilized depth map videos to estimate head movements, which enhanced pain recognition performance when combined with facial expressions.
@@ II Related work
Various compelling methods have been introduced, utilizing video modalities, ranging from hand-crafted feature engineering for facial expressions analysis to employing raw videos with sophisticated deep learning architectures. Werner et al. [41] analyzed facial expressions by extracting features based on point distances and head poses, which were estimated using depth information. By integrating these into a unified feature vector and applying a random forest classifier, they achieved a $76.60\%$ accuracy rate in the binary pain classification task. Patania et al. [50] calculated fiducial points and applied graph neural networks (GNN) to achieve $73.20\%$ accuracy, while the authors in [51] employed vision transformers on raw videos to reach an accuracy of $77.10\%$. Other researchers have made efforts to compute biosignal-related information directly from videos, aiming to enhance assessment accuracy while minimizing the reliance on contact sensors. Yang et al. [52] leveraged raw videos and three-dimensional convolutional neural networks (3D CNNs) to extract facial features and estimate remote photoplethysmography, effectively combining behavioral and physiological cues for enhanced performance. Similarly, the authors in [53] used 3D CNNs to extract pseudo-heart rate from videos, achieving over a $10\%$ increase in accuracy by integrating facial features with heart rate data compared to using videos in isolation.
@@ III Methodology
This section describes the architecture of the proposed framework’s components. It details the foundation model’s multi-task learning-based pre-training, augmentation techniques, and training settings for pre-training and pain assessment tasks. Additionally, it explains the generation of synthetic thermal videos, depth estimation for creating depth videos, and visualization of biosignal modalities used in this study.
@@ III-A1 PainFormer
Before discussing the specific components, we provide preliminary information to outline the general concept of the PainFormer architecture. Every 2D input image $I$ is divided into $n$ non-overlapping patches, with each patch $\in\mathbb{R}^{n\times h\times w\times 3}$, where $h$ and $w$ define the resolution of each patch and are equal to $16\times 16$, and the $3$ indicates the number of color channels. Each patch is linearly projected into a token with dimension $d=768$, followed by a positional encoding layer. Applying Discrete Fourier Transform (DFT), to a 1D input sequence of $N$ elements, $x[n]$, where $n$ ranges from 0 to N1N-1, results in the following expression:
$X[k]=\sum_{n=0}^{N-1}x[n]\,e^{-i2\pi\frac{k}{N}n}=\sum_{n=0}^{N-1}x[n]\,W_{N}^{kn},$ (1)
[…]
@@
where $x[n]$ is the initial time-domain sequence. Furthermore, Eq. (1) can be adapted for two-dimensional inputs, $x[m,n]$, with $0\leq m\leq M-1$ and $0\leq n\leq N-1$, characterized by:
@@
where GELU refers to the Gaussian Error Linear Unit activation function, and DWConv denotes a depthwise convolution layer. In addition, layer normalization is employed before and after the FFT and IFFT processes, refer to Fig. 1(b).
@@ III-A3 Video-Encoder
The architecture of the specific module resembles the Embedding-Mixer; however, for efficiency, it comprises just $1$ layer featuring a $1$ cross-attention module with $1$ head. The number of latent variables, $n$, is set to $256$, while the output embedding length is $40$. This module is employed only within the particular framework for one of the multimodal approaches presented, where video embeddings are integrated with GSR embeddings. The module is illustrated in Fig. 1(e). Note that the Spectral Layer, Self-Attention Layer, Embedding-Mixer, and Video-Encoder all incorporate MLP layers, as depicted in Fig. 1(f, g, h).
@@ III-C Biosignal Visualization
Since the foundation model presented in this study is vision-based, a 2D representation of physiological modalities needs to be employed – representing signals as images has also been successfully demonstrated in other affective-related studies [66, 67, 68]. Four different visualizations are explored: (1) waveform diagrams, which depict the shape and form of a signal as it progresses over time, representing amplitude, frequency, and phase; (2) spectrogram-angle, which illustrates the phase angles of the frequencies; (3) spectrogram-phase, which shows phase information with unwrapping to address discontinuities; and (4) spectrogram-PSD, which represents the power spectral density, exhibiting how power is distributed across frequencies over time. Fig. 3 presents an example of each of these four visualizations.
@@ III-D Foundation Training
TABLE III: Datasets utilized for the multi-task learning-based pre-training process of the PainFormer.
@@ III-F1 BioVid Heat Pain Database
The particular dataset is widely known and established in the field of pain research. It includes facial videos, electrocardiograms, electromyograms, and galvanic skin response measurements from eighty-seven $(n\text{=}87)$ healthy individuals ($44$ males and $43$ females, aged between $20$ and $65$). The experimental setup involved using a thermode to induce pain in the participants’ right arm. Before data collection, each participant’s pain and tolerance thresholds were determined. These thresholds marked the minimum and maximum levels of pain, with two additional intermediate levels, leading to a total of five distinct pain intensities: No Pain (NP), Mild Pain (P1), Moderate Pain (P2), Severe Pain (P3), and Very Severe Pain (P4). The temperature settings used for inducing pain varied within the range from P1 to P4 but never exceeded $50.5^{\circ}\text{C}$. Each participant experienced $20$ pain inductions at each of the four predetermined levels of intensity (P1 to P4). Each stimulus lasted $4s$, followed by a recovery period of $8$ to $12s$. Additionally, $20$ baseline measurements were taken at $32^{\circ}\text{C}$ (NP), resulting in $100$ stimulations per participant, delivered randomly. The data was then processed to create segments of $5.5s$ starting $1s$ after reaching the target temperature for each stimulation. This processing yielded $8,700$ samples, each lasting $5.5s$, distributed evenly across the five pain intensity classes for each modality, covering all $87$ subjects. The videos are recorded at a frame rate of $25$ frames per second (FPS), while the biosignal (ECG, EMG, GSR) recordings are sampled at a rate of 51512 Hz. It should be noted that data from Part A of the dataset were used, where the biosignals had been pre-filtered according to [41].
@@ III-F2 AI4Pain Dataset
[…]